# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v15

@@ 2.7 Asymptotic Behavior for Bound Moment
Formula (2.38) is a special case of Formula (2.37) for symmetric probability distribution. Empirically, Formula (2.38) is confirmed by Formula (2.40), (2.42), and (2.43), and Formula (2.37) is confirmed by Formula (2.45) and (2.51).
$\kappa\ll n$ $:\;\;\;\zeta(n,\kappa)\simeq\rho(\kappa)\frac{\kappa^{n}}{n};$ (2.37)
@@ 2.7.1 Gaussian Input Uncertainty
$\zeta(2n,\kappa)$ $=\int_{-\kappa}^{+\kappa}\tilde{x}^{2n}N(\tilde{x})d\tilde{x}=2N(\kappa)\kappa^{2n}\sum_{j=1}^{\infty}\kappa^{2j-1}\frac{(2n-1)!!}{(2n-1+2j)!!}$ (2.40)
$1<\kappa^{2}\ll 2n:\;\;\;\zeta(2n,\kappa)=2N(\kappa)\frac{\kappa^{2n+1}}{2n+1};$ (2.41)
@@ 2.7.2 Uniform Input Uncertainty
$\zeta(2n)=\int_{-\kappa}^{+\kappa}\tilde{x}^{2n}\frac{1}{2\sqrt{3}}d\tilde{x}=2\frac{1}{2\sqrt{3}}\frac{\kappa^{2n+1}}{2n+1},\;\;0\leq\kappa\leq\sqrt{3};$ (2.42)
2.7.3 Laplace Input Uncertainty
For Laplace distribution [4], Formula (2.43) provides $\zeta(2n,\kappa)$ while $\zeta(2n+1,\kappa)=0$, and Formula (2.44) demonstrates its asymptotic behavior.
$\zeta(2n,\kappa)$ $=\int_{-\kappa}^{+\kappa}\tilde{x}^{2n}\frac{1}{2}e^{-|\tilde{x}|}d\tilde{x}=e^{-\kappa}\kappa^{2n}\sum_{\kappa=1}^{\infty}\kappa^{j}\frac{(2n)!}{(2n+j)!};$ (2.43)
$1<\kappa^{2}\ll 2n:\;\;\;\zeta(2n,\kappa)=2\frac{1}{2}e^{-\kappa}\frac{\kappa^{2n+1}}{2n+1};$ (2.44)
2.7.4 Exponential Input Uncertainty
An exponential distribution [4] is no longer symmetric. Formula (2.45) provides $\zeta(2n,\kappa)$, Formula (2.46) shows the equation for the mean preserving condition, and Formula (2.47) demonstrates its asymptotic behavior.
$\zeta(n,\kappa)$ $=\int_{\varrho}^{\kappa}(\tilde{x}-1)^{n}e^{-\tilde{x}}d\tilde{x}$ (2.45)
$=e^{-\kappa}(\kappa-1)^{n}\sum_{j=1}^{\infty}(\kappa-1)^{j}\frac{n!}{(n+j)!}-e^{-\varrho}(\varrho-1)^{n}\sum_{j=1}^{\infty}(\varrho-1)^{j}\frac{n!}{(n+j)!};$
$\varrho e^{-\varrho}=\kappa e^{-\kappa}\Longrightarrow 0<\varrho<1<\kappa;$ (2.46)
$1<\kappa\ll n:\;\;\;\zeta(n,\kappa)=e^{-\kappa}\frac{\kappa^{n+1}}{n+1};$ (2.47)
2.7.5 Simplified Gamma Input Uncertainty
The gamma distribution [4] $\rho(\tilde{x})=x^{\alpha-1}e^{-x/\lambda}\lambda^{\alpha}/\Gamma(\alpha),\tilde{x},\alpha,\lambda>0$ is simplified as Formula (2.48), whose mean and deviation are $m+1$ and $\sqrt{m+1}$, respectively. Formula (2.49) provides $\zeta(2n,\kappa)$, Formula (2.50) shows the equation for the mean preserving condition, and Formula (2.51) demonstrates its asymptotic behavior.
[…]
$\zeta(n,\kappa)$ $=\int_{\varrho}^{\kappa}\left(\frac{\tilde{x}-(m+1)}{\sqrt{m+1}}\right)^{n}\rho(\tilde{x})d\tilde{x}$ (2.49)
$e^{-\varrho}\varrho^{m+1}=e^{-\kappa}\kappa^{m+1};\Longrightarrow 0<\varrho<m+1<\kappa;$ (2.50)
$1<\kappa\ll n:\;\;\;\zeta(n,\kappa)=\frac{1}{m!}e^{-\kappa}\frac{\kappa^{m+n+1}}{m+n+1};$ (2.51)
@@ 2.8 Convergence of Variance
When $n\rightarrow\infty$, $\zeta(j,\kappa)\zeta(n-j,\kappa)$ will be ignored in Formula (2.3) because $\zeta(j,\kappa)\zeta(n-j,\kappa)\ll\zeta(n,\kappa)$. The asymptotic behavior of $\zeta(n,\kappa)$ when $n\rightarrow\infty$ determines if Formula (2.3) will converge.
@@ 2.9 Statistical Bounding
When the underlying distribution is Normal, Formula (2.54) and (2.55) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Normal error function [4]. Figure 6 shows that $\epsilon(\kappa_{s})<\epsilon(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\epsilon(\kappa_{s},N)=\epsilon(\kappa_{s})$. It also shows that $\kappa(\kappa_{s},N)<\kappa_{s}$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$. Figure 7 further demonstrates that for smaller $\kappa_{s}$, $\kappa(\kappa_{s},N)$ approaches $\kappa_{s}$ more rapidly as $N$ increases (e.g., $\kappa(2,100)\simeq 2$ vs $\kappa(5,1000)\simeq 5$), but converges to a larger stable bounding leakage (e.g., $\epsilon(2)=4.55\;10^{-2}$ vs $\epsilon(5)=5.73\;10^{-7}$). Figure 7 also indicates that when $N\geq 30$, the difference between $\epsilon(4,N)$ and $\epsilon(5,N)$ is less than $10^{-3}$, suggesting that $\kappa(\kappa_{s},N)$ becomes stable when $\kappa_{s}\geq 4$. Moreover, according to the 5-$\sigma$ rule, $\kappa(5,N)$ in Figure 6 should be used as $\kappa$ in Formula (2.7).
$\kappa_{s}=\sqrt{3}$ and $\kappa_{s}=5$ are defined as the ideal bounding ranges $\hat{\kappa}$ for the uniform and Gaussian distributions respectively.
@@ 2.10 Ideal Statistics
Define the stable variance when the sample count $N$ is sufficient large as the ideal variance $\widehat{\delta^{2}}f$, which assumes that the sample count $N$ is effectively infinite for each input. Ideal variances are consistent with traditional variances [4].
@@
Statistical Taylor expansion outputs ideal uncertainty, ideal uncertainty bias, and ideal ratio for the chosen ideal bounding ranges, in which the ideal ratio quantifies the reliability of the ideal uncertainty and ideal uncertainty bias.
@@ 2.11 Dependency Tracing
$\delta^{2}\left(\frac{f^{(m)}_{x}\tilde{x}^{m}}{m!}+\frac{f^{(n)}_{x}\tilde{x}^{n}}{n!}\right)=$ $\;\sigma^{2m}(\frac{f^{(m)}_{x}}{m!})^{2}\eta(2m)+\sigma^{2n}(\frac{f^{(n)}_{x}}{n!})^{2}\eta(2n)$ (2.62)
$+2\sigma^{m+n}\left(\frac{f^{(m)}_{x}f^{(n)}_{x}}{m!\;n!}\eta(m+n)-\frac{f^{(m)}_{x}}{m!}\eta(m)\;\frac{f^{(n)}_{x}}{n!}\eta(n)\right);$
@@
  • Formula (2.58) expresses $\delta^{2}(f+g)$, whose dependency tracing is illustrated by $\delta^{2}(f-f)=0$, and $\delta^{2}(f(x)+g(y))=\delta^{2}f+\delta^{2}g$, with the latter corresponding to Formula (2.15). Formula (2.62) shows that Formula (2.3) applies Formula (2.58) between any two terms in the Taylor expansion in Formula (2.8).
@@ 9.10 Numerical Error Resonance Using Library Sine
In the reverse transformation, value errors exhibit a clear trend of increasing with the time index, as shown in Figure 51. These large value errors appear systematic rather than random and visually resemble a resonant pattern. Similar increases are observed at in other frequencies and FFT orders, as well as in computational results obtained using mathematical libraries such as SciPy. For example, Figure LABEL:fig:_FFT_Cost_Clean_6_6_Wave_Lib shows that the FFT value error waveform of $\cos(6\frac{2\pi}{2^{6}}j)$ has much stronger resonant pattern than that of $\sin(3\frac{2\pi}{2^{6}}j)$ in Figure 46. In the forward transformation, although there is no visible resonant pattern in FFT value error specturm, the intensity of value errors increase with FFT order. In contrast, such resonance is absent for the roundtrip transformation as shown in Figure 50. Also, it is completely absent when using the Quart sine function, as seen in Figure 39 and 40. Figure 51 and Figure 52 demonstrate that the error deviations increase with sine or cosine frequency. Figure 35 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity of an integer multiply of $\pi$, producing the resonant pattern in Figure 45, 46, and 49. At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger. To suppress this numerical error resonances, an input noise level of approximately $10^{-14}$ input noise must be added to the sine or cosine signals.
@@ 11.3 Improvements Needed
Bound momentum $\zeta(2n)$ needs to be extended to all probability distributions. The necessary and sufficient conditions for Formula (2.37) need to be found, to extend the applicable types of distributions for the statistical Taylor expansion from the most common ones as enumerated in Section 2. The statistical bounding to find the bounding range $\kappa$ for the bounded moment needs to be extended to discrete distributions.
The choice of ideal bounding range $\hat{\kappa}$ needs further study. For example, the choice of $\hat{\kappa}=5$ for Normal distribution is based on the $5-\sigma$ rule in practice only. Such choice certainly can not be applied to other distributions. Also, the choice of ideal bounding range $\hat{\kappa}$ determines the convergence of both $\log(x)$ and $x^{c}$ as $P(x)\lesssim 1/\hat{\kappa}$, which needs better mathematical understanding.

### v16

@@ 2.7 Asymptotic Behavior for Bound Moment
When $\rho(\tilde{z})>0$ and $\frac{d\rho(\tilde{z})}{d\tilde{z}}$ is bounded in $\tilde{z}\in[0,\kappa]$, Formula (2.38) and (2.39) leads to Formula (2.40). Formula (2.37) and (2.40) conclude Formula (2.41) and (2.42).
$\int_{0}^{\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}$ $=\rho(\kappa)\frac{\kappa^{n+1}}{n+1}-\frac{1}{n+1}\int_{0}^{\kappa}\tilde{z}^{n+1}\frac{d\rho(\tilde{z})}{d\tilde{z}}d\tilde{z}$; (2.37)
@@
[…]
$\lim_{n\rightarrow+\infty}$ $\int_{0}^{\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}=\rho(\kappa)\frac{\kappa^{n+1}}{n+1};$ (2.40)
$|\varrho|=\kappa$ $:\;\;\;\lim_{n\rightarrow+\infty}\zeta(2n,\kappa)=2\rho(\kappa)\frac{\kappa^{2n+1}}{2n+1};$ (2.41)
$|\varrho|<\kappa$ $:\;\;\;\lim_{n\rightarrow+\infty}\zeta(n,\kappa)=\rho(\varrho)\frac{\varrho^{n+1}}{n+1}+\rho(\kappa)\frac{\kappa^{n+1}}{n+1}\simeq\rho(\kappa)\frac{\kappa^{n+1}}{n+1};$ (2.42)
@@ 2.7.1 Gaussian Input Uncertainty
$\zeta(2n,\kappa)$ $=\int_{-\kappa}^{+\kappa}\tilde{z}^{2n}N(\tilde{z})d\tilde{z}=2N(\kappa)\kappa^{2n}\sum_{j=1}^{\infty}\kappa^{2j-1}\frac{(2n-1)!!}{(2n-1+2j)!!}$ (2.44)
$\lim_{n\rightarrow+\infty}$ $\zeta(2n,\kappa)=2N(\kappa)\frac{\kappa^{2n+1}}{2n+1};$ (2.45)
@@ 2.7.2 Uniform Input Uncertainty
$\zeta(2n)=\int_{-\kappa}^{+\kappa}\tilde{z}^{2n}\frac{1}{2\sqrt{3}}d\tilde{z}=2\rho(\kappa)\frac{\kappa^{2n+1}}{2n+1},\;\;0\leq\kappa\leq\sqrt{3};$ (2.46)
@@ 2.8 Convergence of Variance
$\lim_{n\rightarrow+\infty}$ $\zeta(n)-\zeta(j)\zeta(n-j)=\frac{\kappa^{n}}{n};$ (2.47)
For input distribution without bounding, Formula (2.3) may not converge, e.g., when the input uncertainty is Laplace [4] so that $\zeta(2n)=(2n)!$. Thus, it is necessary that the input uncertainty is bounded. Formula (2.47) shows that the asymptotic behavior of $\zeta(n,\kappa)$ when $n\rightarrow\infty$ determines if Formula (2.3) will converge.
@@ 2.9 Statistical Bounding
When the underlying distribution is Normal, Formula (2.50) and (2.51) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Normal error function [4]. Figure 6 shows that $\epsilon(\kappa_{s})<\epsilon(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\epsilon(\kappa_{s},N)=\epsilon(\kappa_{s})$ along the y-axis direction. It also shows that $\kappa(\kappa_{s},N)<\kappa_{s}$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$ along the x-axis direction. Figure 7 further demonstrates that for smaller $\kappa_{s}$, $\kappa(\kappa_{s},N)$ approaches $\kappa_{s}$ more rapidly as $N$ increases (e.g., $\kappa(2,100)\simeq 2$ vs $\kappa(5,1000)\simeq 5$), but converges to a larger stable bounding leakage (e.g., $\epsilon(2)=4.55\;10^{-2}$ vs $\epsilon(5)=5.73\;10^{-7}$). Figure 7 also indicates that when $N\geq 30$, the difference between $\epsilon(4,N)$ and $\epsilon(5,N)$ is less than $10^{-3}$, suggesting that $\kappa(\kappa_{s},N)$ becomes stable when $\kappa_{s}\geq 4$. Moreover, according to the 5-$\sigma$ rule, $\kappa(5,N)$ in Figure 6 should be used as $\kappa$ in Formula (2.7).
[…]
@@ 2.10 Ideal Statistics
Define the stable variance when the sample count $N$ is sufficient large as the ideal variance $\widehat{\delta^{2}}f$, which assumes that the sample count $N$ is effectively infinite for each input.
@@
In statistical Taylor expansion, every calculation should be carried out twice:
  1. 1.
[…]
  2. 2.
    With the real $\zeta(n,\kappa)$ of each input given the sample count and input uncertainty distribution, to obtain the ideal ratio to quantify the reliability of the ideal variance and ideal uncertainty bias as a value in the range of $[0,1)$.
@@ 2.11 Dependency Tracing
  • Formula (2.54) expresses $\delta^{2}(f+g)$, whose dependency tracing is illustrated by $\delta^{2}(f-f)=0$, and $\delta^{2}(f(x)+g(y))=\delta^{2}f+\delta^{2}g$, with the latter corresponding to Formula (2.15). Formula (LABEL:eqn:_Taylor_inner_dependency) shows that Formula (2.3) applies Formula (2.54) between any two terms in the Taylor expansion in Formula (2.8).
@@ 9.10 Numerical Error Resonance Using Library Sine
In the reverse transformation, value errors exhibit a clear trend of increasing with the time index, as shown in Figure 51. These large value errors appear systematic rather than random and visually resemble a resonant pattern. Similar increases are observed at in other frequencies and FFT orders, as well as in computational results obtained using mathematical libraries such as SciPy. For example, Figure 49 shows that the FFT value error waveform of $\cos(6\frac{2\pi}{2^{6}}j)$ has much stronger resonant pattern than that of $\sin(3\frac{2\pi}{2^{6}}j)$ in Figure 46. In the forward transformation, although there is no visible resonant pattern in FFT value error specturm, the intensity of value errors increase with FFT order. In contrast, such resonance is absent for the roundtrip transformation as shown in Figure 50. Also, it is completely absent when using the Quart sine function, as seen in Figure 39 and 40. Figure 51 and Figure 52 demonstrate that the error deviations increase with sine or cosine frequency. Figure 35 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity of an integer multiply of $\pi$, producing the resonant pattern in Figure 45, 46, and 49. At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger. To suppress this numerical error resonances, an input noise level of approximately $10^{-14}$ input noise must be added to the sine or cosine signals.
@@ 11.3 Improvements Needed
Bound momentum $\zeta(2n)$ needs to be extended to all probability distributions. The statistical bounding to find the bounding range $\kappa$ for the bounded moment needs to be extended to discrete distributions.
The choice of ideal bounding range $\hat{\kappa}$ needs further study. For example, the choice of $\hat{\kappa}=5$ for Normal distribution is based on the $5-\sigma$ rule in practice only. Such choice certainly can not be applied to other distributions. Also, the choice of ideal bounding range $\hat{\kappa}$ determines the convergence of both $\log(x)$ and $x^{c}$ as $P(x)\lesssim 1/\hat{\kappa}$, which needs better mathematical understanding. The condition for Formula (2.47) seems to be as generic as the definition in Formula (2.7) which needs further proof.